### v1

Abstract
Large Language Models (LLMs), such as the GPT-4 and LLaMA families, have demonstrated considerable success across diverse tasks, including multiple-choice questions (MCQs). However, these models exhibit a positional bias, particularly an even worse “anchored bias” in the GPT-2 family, where they consistently favour the first choice ‘A’ in MCQs during inference. This anchored bias challenges the integrity of GPT-2’s decision-making process, as it skews performance based on the position rather than the content of the choices in MCQs. In this study, we utilise the mechanistic interpretability approach to identify the internal modules within GPT-2 models responsible for this bias. We focus on the Multi-Layer Perceptron (MLP) layers and attention heads, using the “logit lens” method to trace and modify the specific value vectors that contribute to the bias. By updating these vectors within MLP and recalibrating attention patterns to neutralise the preference for the first choice ‘A’, we effectively mitigate the anchored bias. Our interventions not only correct the bias but also improve the overall MCQ prediction accuracy for the GPT-2 family across various datasets. This work represents the first comprehensive mechanistic analysis of anchored bias in MCQs within the GPT-2 models, introducing targeted, minimal-intervention strategies that significantly enhance GPT2 model robustness and accuracy in MCQs. Our code is available at https://github.com/ruizheliUOA/Anchored_Bias_GPT2.
1 Introduction
Large Language Models (LLMs) exhibit remarkable capabilities across a wide array of tasks, including multiple-choice question (MCQ) (Robinson and Wingate,, 2022), which are largely attributed to the advancements in the Transformer backbone. These models not only excel at reasoning but also demonstrate significant inductive capabilities, which make them highly effective in different domains (Chen et al.,, 2023; Team et al.,, 2023; Anil et al.,, 2023).
Despite their success, recent studies have uncovered a notable flaw: LLMs exhibit a positional bias when tasked with MCQs. Specifically, the performance of these models (e.g., LLaMA (Touvron et al., 2023a, ), LLaMA2 (Touvron et al., 2023b, ), GPT-4 (Achiam et al.,, 2023)) varies significantly depending on the position of the correct answer within the given choices (Pezeshkpour and Hruschka,, 2023; Zheng et al., 2024a, ). We further observe that this vulnerability to positional bias is even worse in the GPT-2 family, ranging from GPT2-Small-124M to GPT2-XL-1.5B (Radford et al.,, 2019). Our investigations reveal that GPT-2 models consistently favour the first choice ‘A’, regardless of the actual position in the input prompt where the correct answer choice is placed, which we term as “anchored bias”.
Previous work primarily mitigated positional bias in MCQ by analysing the impact of different prompt structures (Pezeshkpour and Hruschka,, 2023) or by estimating different datasets’ prior bias based on test samples (Zheng et al., 2024a, ). Such approaches often remain superficial, merely altering the prompt presentation, or are hard to generalise when there is not a large test dataset. Remarkably, there has been a lack of investigation into the internal mechanisms of LLMs that contribute to the anchored bias and strategies to mitigate it without the need for prompt engineering or prior estimation.
Figure 1: MCQ prompt paradigm used in GPT2-Small and next token logit rankings with probability during inference. Regardless of the order in which correct answer choices are placed in the prompt except ‘A’, GPT2-Small always give a higher logit score to the choice immediately following the Answer Choices:, i.e., A, where Refer to caption represents the anchored bias for the incorrect choices (the correct choices should be B and C for this example).
We apply mechanistic interpretability to reverse-engineer the internal workings of the GPT-2 family to understand the origins and extent of the anchored bias. We quantitatively demonstrate that the GPT-2 Small, Medium, Large, and XL models exhibit this anchored bias with significant regularity across various MCQ datasets, ranging from 2 choices to 5 choices settings. Our detailed analysis using the “logit lens” (Nostalgebraist,, 2020) approach localises Multi-Layer Perceptron (MLP) layers with specific dimensionality and attention heads that disproportionately influence this anchored bias. We find that certain value vectors in the MLP, which inherently harbour this bias, and specific attention heads pay more weight on the ‘A’ position over the correct answer choice positions in the input prompt.
Inspired from (Geva et al.,, 2021, 2022) where MLPs can be treated as key-value memories, we use a straightforward yet potent method (Dai et al.,, 2022) to update these critical value vectors in the MLP, effectively mitigate the anchored bias. This adjustment not only mitigates the anchored bias but also enhances the overall MCQ prediction accuracy of the GPT-2 family over 70% averaged across various MCQ datasets and all GPT-2 family. Additionally, we propose a novel strategy to recalibrate the attention patterns by swapping the attention weight between the anchored position and the correct answer choice position. This strategy also mitigates the anchored bias to a certain extent, especially for the classification accuracy improvement of the Indirect Obejct Identification (IOI) dataset (Wang et al.,, 2022) over 90% on GPT2-Medium.
In conclusion, to our best knowledge, this work is the first comprehensive mechanistic analysis of the intrinsic anchored bias in MCQ tasks across the entire GPT-2 family. By identifying and rectifying the critical value vectors within MLP and attention heads responsible for this bias, we introduce novel, minimal-intervention strategies that significantly reduce GPT-2 models’ vulnerability and enhance robustness against anchored bias in MCQ task.
2 Background: Large Language Models and Mechanistic Interpretability
Figure 2: Left: the architecture of LLMs, where the input prompt is encoded via $W_{E}$, and then the processed information via attention and MLP layer is accumulated back to the residual stream $\mathbf{X}_{\text{mid}}^{\ell}$ and $\mathbf{X}^{\ell}$ at layer $\ell$, respectively. Finally, the residual stream at $L$ layer is unembedded as logits and normalised as a probability distribution for next token prediction. Right: logit lens (Nostalgebraist,, 2020) is used to investigate the contribution of attention pattern and MLP module for the next token prediction.
Architecture of LLMs.
We focus on the autoregressive Transformer-based LLM architecture (Vaswani et al.,, 2017) based on prior works (Geva et al.,, 2021; Elhage et al.,, 2021; Geva et al.,, 2022; Dai et al.,, 2022; Meng et al., 2022a, ; Meng et al., 2022b, ; Yuksekgonul et al.,, 2024) with simplifications in certain explanations. Given an input prompt containing $T$ tokens $(t_{1},\ldots,t_{T})$ and each token $t_{i}$ belonging to a vocabulary $\mathcal{V}$, tokens are initially encoded by $d$-dimensional vectors $\mathbf{x}_{i}^{0}\in\mathbb{R}^{d}$ using an embedding matrix $W_{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$.
The architecture has $L$ layers, and each layer consists of attention and MLP modules, which transform token embeddings to residual streams $(\mathbf{x}_{1}^{\ell},\ldots,\mathbf{x}_{T}^{\ell})\in\mathbf{X}^{\ell}$ at layer $\ell$, where $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$. The residual stream at layer $\ell$ is a place where all attention and MLP modules at layer $\ell$ read from and write to (Elhage et al.,, 2021), and it is updated by the following equation for token $i$ at layer $\ell$:
$\mathbf{x}_{i}^{\ell}=\mathbf{x}_{i}^{\ell-1}+\mathbf{a}_{i}^{\ell}+\mathbf{m}_{i}^{\ell}$ (1)
Here, $\mathbf{a}_{i}^{\ell}$ is the attention contribution for token $i$ and $\mathbf{m}_{i}^{\ell}$ is the MLP contribution at layer $\ell$ [^1] We omit the layer normalisation of attention and MLP modules at layer $\ell$ for simplification.. At $L$ layer, the predicted probability distribution for the next token $\mathcal{P}(t_{T+1}|t_{1:T})$ is produced following:
$\mathcal{P}(t_{T+1}|t_{1:T})=\text{Softmax}\left(W_{U}\sigma(\mathbf{x}_{T}^{L})\right)$ (2)
where $W_{U}\in\mathbb{R}^{d\times|\mathcal{V}|}$ is unembedding matrix, $\sigma(\cdot)$ is pre-unembedding layer normalisation.
The attention module mainly updates each token residual stream $\mathbf{x}_{i}^{l-1}$ by attending to all previous tokens in parallel. Specifically, the attention module contains $QK$ and $OV$ circuits, where the former operates $W_{Q},W_{K}\in\mathbb{R}^{d\times d}$ matrices and the latter operates $W_{O},W_{V}\in\mathbb{R}^{d\times d}$ matrices, respectively. Normally, $QK$ circuit determines the attention pattern $A^{\ell}$, i.e., where information is moved to and from the residual stream. $OV$ circuit further determines the attention output $\mathbf{a}_{i}^{\ell}$ based on the fixed attention pattern, i.e., what information is from the previous tokens’ position to the current token position (Elhage et al.,, 2021):
$A^{\ell,h}=\text{Softmax}\left(\frac{(\mathbf{X}^{\ell-1}W_{Q}^{\ell,h})(\mathbf{X}^{\ell-1}W_{K}^{\ell,h})^{T}}{\sqrt{d_{h}}}\right)$ (3)
$\mathbf{a}_{i,j}^{\ell}=\sum_{h=1}^{H}A_{i,j}^{\ell,h}(\mathbf{x}_{j}^{\ell-1}W_{V}^{\ell,h})W_{O}^{\ell,h}=\sum_{h=1}^{H}\mathbf{r}_{i,j}^{\ell,h}$ (4)
where $\mathbf{a}_{i,j}^{\ell}$ indicates the attention contribution from token $i$ to token $j$, and $\mathbf{a}_{i}^{\ell}=\sum_{j=1}^{T}\mathbf{a}_{i,j}^{\ell}$. Attention pattern $A^{\ell,h}\in\mathbb{R}^{T\times T}$ is a lower triangular weight matrix calculated by the $h$-th attention head at layer $\ell$, representing that each token can only attend to previous tokens within autoregressive Transformer-based LLMs. All matrices are split into multiple attention heads, i.e., $W_{Q}^{\ell,h},W_{K}^{\ell,h},W_{V}^{\ell,h}\in\mathbb{R}^{d\times d_{h}}$, and $W_{O}^{\ell,h}\in\mathbb{R}^{d_{h}\times d}$ for head $h$. $d_{h}$ is the dimensionality of each head, $H$ represents the total number of attention heads, and $d_{h}=d/H$. $A_{i,j}^{\ell,h}$ is the $i$-th row and $j$-th column entry of $A^{\ell,h}$, and $\mathbf{r}_{i,j}^{l,h}$ indicates the weighted average values where token $i$ attend to token $j$ by head $h$ at the layer $\ell$, and $\mathbf{r}_{i}^{\ell,h}=\sum_{j=1}^{T}\mathbf{r}_{i,j}^{\ell,h}$.
For MLP module, it receives the $\mathbf{x}_{i}^{\ell-1}$ as input and updates following:
$\mathbf{m}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})W_{\text{out}}^{\ell}$ (5)
where $\gamma(\cdot)$ is activation function, $W_{\text{in}}^{\ell}\in\mathbb{R}^{d\times d_{\text{m}}}$, and $W_{\text{out}}^{\ell}\in\mathbb{R}^{d_{\text{m}}\times d}$. $d_{\text{m}}$ is the dimensionality of MLP module, which is larger than $d$. MLP module is normally treated as key-value memories (Geva et al.,, 2021; Elhage et al.,, 2021; Geva et al.,, 2022; Dai et al.,, 2022), where columns of $W_{\text{in}[:,i]}^{\ell}$ and rows of $W_{\text{out}[i,:]}^{\ell}$ are viewed as keys and values, respectively. Given the input $\mathbf{x}_{i}^{\ell-1}$, the keys of MLP produce a vector of cofficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, and they weights the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$. Therefore, Eq. 5 can be reformatted as [^2] We omit all bias $b_{Q},b_{K},b_{O},b_{V},b_{\text{in}},b_{\text{out}},b_{\text{U}}$ for simplification.:
$\mathbf{m}_{i}^{\ell}=\sum_{n=1}^{d_{\text{m}}}\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n}$ (6)
Logit lens.
Logit lens is a mechanistic interpretability approach to investigate the contribution of the intermediate layer representation in the autoregressive Transformer-based LLMs (Nostalgebraist,, 2020). Based on the architecture of LLMs above, the $\mathcal{P}(t_{T+1}|t_{1:T})$ at layer $L$ is the production of linear softmax of logits unembedded via $W_{U}$, which is the sum of input $\mathbf{x}_{i}^{0}$ and attention and MLP contributions at each layer $\ell$. Therefore, logit lens can be used to measure the weighted attention value of each head $\mathbf{r}_{i}^{\ell,h}\in\mathbb{R}^{d}$, each weighted value vector $\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n}\in\mathbb{R}^{d}$ at $n$-th dimensionality in MLP and intermediate residual stream $\mathbf{x}_{i}^{\ell}\in\mathbb{R}^{d}$ for token $i$:
$\text{logit}_{i}^{\ell,h}(\mathbf{r}_{i}^{\ell,h})=W_{U}\sigma(\mathbf{r}_{i}^{\ell,h})\quad\text{logit}_{i}^{\ell,n}(\mathbf{m}_{i}^{\ell,n})=W_{U}\sigma(\mathbf{k}_{i}^{\ell,n}\mathbf{v}_{i}^{\ell,n})\quad\text{logit}_{i}^{\ell}(\mathbf{x}_{i}^{\ell})=W_{U}\sigma(\mathbf{x}_{i}^{\ell})$ (7)
3 Preliminaries: Zero-shot Learning with MCQs
Zero-shot learning.
We mainly focus on the zero-shot learning regarding each GPT2 model, i.e., the input prompt is formatted as “Question: <Question sample> Answer Choices: <Multiple Choices> Answer:”, which is explained in Fig. 1. After encoding the input prompt, GPT2 model will decode the next token prediction, which is expected as the correct answer choice.
Table 1: The distribution of correct options on each training dataset. IOI, LD, and Greater-than datasets are manual-synthesised and we did not choose or place the correct option at A. For test datasets, we only select samples whose correct options are not ‘A’ to avoid overlap between anchored predictions from GPT2 models and the correct option. ()($\cdot$)( ⋅ ) indicates the number of choices.
Datasets Train Test A (%) B (%) C (%) D (%) E (%)
IOI (2) - 1000 0 100 - - -
LD (3) - 200 0 50 50 - -
Greater (4) - 1000 0 33.33 33.33 $3$
ARC (4) 1.12k 907 20.82 26.18 25.65 25.29 -
CSQA (5) 9.74k 982 19.60 20.25 19.98 20.38 19.79
Datasets and models.
To comprehensively verify and evaluate the anchored bias of GPT2 family, we consider 5 datasets, which include different numbers of choices from 2 to 5. Indirect Object Identification (IOI) Wang et al., (2022): IOI is a manually synthesised corpus used to understand a specific natural language task, where sentences such as “Afterwards Lisa and Rachel went to the garden, and Lisa gave a bone to” should be followed with “Rachel”. When two names are included in such sentences, the predicted name should not be the subject of the last clause. This task has been verified that GPT2 family works well (Wang et al.,, 2022; Merullo et al.,, 2024). However, we found that GPT2 family immediately fails this task if the input prompt is formatted as MCQ, where the incorrect subject of the last clause is placed in the ‘A’ choice. Greater-than task (Greater) Hanna et al., (2024): Greater-than task is also a manually synthesised corpus, which is used to evaluate GPT2’s ability to sentences such as “The war lasted from the year 1732 to the year 17”, and model will predict valid two-digit end years, i.e., years > 32. However, we found the same anchored bias like IOI when this task is formatted as MCQ in Fig. 1, and GPT2 family also fails to predict valid years and the prediction is always anchored at incorrect choice ‘A’. Logical Deduction of the Big-Bench (LD) [^3] https://github.com/google/BIG-bench/blob/main/bigbench/benchmark_tasks/logical_deduction/three_objects/task.json Srivastava et al., (2023): LD is a subtask which evaluates three-object logical deduction tasks, and it is used to measure whether model can parse information about multiple choices and their mutual relationships. ARC-Challenge (ARC) Clark et al., (2018) and CommensenseQA (CSQA) Talmor et al., (2019) are commonly-used MCQ benchmarks to evaluate LLMs. For models, we comprehensively evaluate the GPT2 family, i.e., GPT2-Small-124M, GPT2-Medium-355M, GPT2-Large-774M, and GPT2-XL-1.5B Radford et al., (2019).
Table 2: The distribution of anchored bias ‘A’ happened for GPT2 family across different datasets.
Dist. (%) GPT2-Small GPT2-Medium GPT2-Large GPT2-XL
IOI (2) 45.5 97.4 100.0 85.8
LD (3) 63.0 94.0 100.0 17.0
Greater (4) 32.1 95.0 99.5 98.0
ARC (4) 54.6 91.6 97.6 69.9
CSQA (5) 34.8 81.5 99.6 97.7
Evaluation metrics.
We use logit lens (Nostalgebraist,, 2020) introduced in § 2 to localise the specific layer of MLP and specific attention head which contribute to the anchored bias (more details in § 4). Moreover, we use MLP contribution (Geva et al.,, 2022) to locate the specific dimensionality from $W_{\text{out}}^{\ell}$ which leads to anchored bias. Regarding mitigating anchored bias, we use classification accuracy to evaluate whether the anchored bias can be mitigated and GPT2 family can successfully predict correct choices in MCQ.
4 Discovering Anchored Bias in MCQs
Frequency of anchored bias in GPT2 family across all datasets.
As introduced in § 3, we use 5 different datasets with choices from 2 to 5 to investigate the anchored bias. Table 1 shows that the correct choices distribution of ARC and CSQA training dataset is balanced from ‘A’ to ‘E’. In addition, IOI, LD and Greater test datasets are manually synthesised, and we did not choose or place the correct choice at ‘A’. For all test datasets, we only select samples whose correct choice is not ‘A’ to avoid introducing extra bias, i.e., the mix-up between correct prediction and anchored bias of GPT2 family. Based on the randomly sampled test datasets in Table 1, we further calculate the distribution of anchored bias ‘A’ happened within each test dataset in Table 2. We can find that GPT2-Large and GPT2-Medium have the most serious anchored bias, and GPT2-XL and GPT2-Small have relatively less serious issues. Based on this situation, we mainly focus on investigating test samples which have anchored bias for each dataset. Table 7 in Appendix shows the number of test samples for each dataset and GPT2 model, where Infer. is used to localise and mitigate anchored bias and Eva. is used to verify the performance of mitigation.
Figure 3: MLP logit difference between anchored bias token ‘A’ and correct tokens (one of B,C,D,E), i.e., $\text{logit}_{T}^{\ell}[\texttt{A}](\mathbf{m}_{T}^{\ell})-\text{logit}_{T}^{\ell}[\texttt{B/C/D/E}](\mathbf{m}_{T}^{\ell})$ which is averaged within GPT-2 family across all layers and all datasets. The deeper the blue blocks are at each layer, the more serious the anchored bias is, and vice versa.
Locating MLP of GPT2 family for anchored bias.
We first investigate MLP modules within GPT2 family for anchored bias. Inspired from Geva et al., (2021, 2022), the MLP modules can be regarded as key-value memories. As introduced in § 2 and Fig. 2, the keys of MLP module is a vector of coefficients $\mathbf{k}_{i}^{\ell}=\gamma(W_{\text{in}}^{\ell}\mathbf{x}_{i}^{\ell-1})\in\mathbb{R}^{d_{\text{m}}}$, which dynamically controls the contributions of the corresponding values $\mathbf{v}_{i}^{\ell}$ in $W_{\text{out}}^{\ell}$ based on different input prompts. The value $\mathbf{v}_{i}^{\ell}$ is treated as a memory bank which stores knowledge after the model pertaining.
Based on the consensus about MLP module, we aim to solve these research questions: 1) Is MLP responsible for the anchored bias in GPT2 family? 2) Which layer and dimensionality of MLP is anchored bias relevant to? 3) Is this bias stored as knowledge in a specific value vector of $W_{\text{out}}^{\ell}$?
We use logit lens (Nostalgebraist,, 2020) to calculate logit of the final input prompt token contributing to incorrect choice token ‘A’ and correct choice token ‘B/C/D/E’ based on different datasets using Eq. 6 and Eq. 7444The reason why we focus on the final input token is that the information inside of autoregressive transformer-based model will accumulate to the final input token for the next token prediction.:
$\text{logit}_{T}^{\ell}(\mathbf{m}_{T}^{\ell})[\texttt{A}]=W_{U}[\texttt{A}]\sigma(\mathbf{m}_{T}^{\ell})\quad\text{logit}_{T}^{\ell}(\mathbf{m}_{T}^{\ell})[\texttt{B/C/D/E}]=W_{U}[\texttt{B/C/D/E}]\sigma(\mathbf{m}_{T}^{\ell})$ (8)
[…]
We calculate the MLP logit difference between anchored bias token ‘A’ and correct choice token B/C/D/E averaged across all layers and datasets for each GPT2 model using Infer. test samples. As shown in Fig. 3, layer 9 in GPT2-Small, layer 20 in GPT2-Medium, layer 34 in GPT2-Large, and layer 37/38/44 in GPT2-XL are dominant layers [^5] In this work, the layer number and head number start from 0. related to anchored bias. In addition, these layers are much closer to the final layer of GPT2, which agrees with Geva et al., (2022); Gurnee et al., (2023)’s finding that higher layers in GPT2 are relevant to semantic concepts or complicated tasks. We also notice that the last one or two layers in each GPT2 model do not have anchored bias at all, and they contribute more logits to correct choice token ‘B/C/D/E’ to ‘A’.
Based on the pattern from Fig. 3, we further use MLP contribution (Geva et al.,, 2022) to localise the specific dimensionality from $W_{\text{out}}^{\ell}$ in these identified layers leading to anchored bias:
$\text{Contrib}(\mathbf{v}_{T}^{\ell,n})=|\mathbf{k}_{T}^{\ell,n}|||\mathbf{v}_{T}^{\ell,n}||$ (9)
where $|\mathbf{k}_{T}^{\ell,n}|$ is the absolute value of the coefficient $\mathbf{k}_{T}^{\ell,n}$, and $n\in d_{m}$. After using Eq. 9, we can locate the top-10 most dominant weighted value vector $\mathbf{k}_{T}^{\ell,n}\mathbf{v}_{T}^{\ell,n}$ and dimensionality with the largest contribution of the final input prompt token. We further calculate logit difference of these identified layers and dominant dimensionality in MLP of the final input token contributing anchored bias token ‘A’ and correct choice tokens B/C/D/E using Eq. 7, i.e., $\text{logit}_{T}^{\ell,n}[\texttt{A}](\mathbf{m}_{T}^{\ell,n})-\text{logit}_{T}^{\ell,n}[\texttt{B/C/D/E}](\mathbf{m}_{T}^{\ell,n})$. Then we select candidates among the top-10 dominant dimensionality where difference score is larger than 4.
Table 3 Vector column demonstrates the specific value vectors $\mathbf{v}^{\ell,n}$ which are responsible for anchored bias. For each value vector, we also calculate the percentage of how frequently it is recognised as an anchored-bias vector across all datasets for each GPT2 model. We can find that most identified value vectors have more than 50% chance with anchored bias happening across all datasets and different GPT2 models. To further verify whether these value vectors store anchored knowledge bias, we unembeded each value vector logit and selected the top-10 tokens with the highest probability. As shown in Table 3, we can find that most top-10 tokens within each value vector are relevant to ‘A’, e.g., ␣A, A, ␣a, a, ␣first, etc. This finding proves that some value vectors in $W_{\text{out}}$ of GPT2 family store knowledge bias after pertaining, and these knowledge biases will become anchored bias when the input prompt is formatted as MCQ. In addition, most unembeded tokens are stopwords, e.g., pronouns, articles, prepositions, etc, which also agrees with the findings that “stopwords/punctuation” are commonly distributed in the value vectors of MLP (Geva et al.,, 2022).
Table 3: Identified anchored-bias value vectors $\mathbf{v}^{\ell,n}$ of $n$-row of $W_{\text{out}}^{\ell}$ at layer $\ell$ for each GPT2 model, where the percentage indicates how frequently the specific $\mathbf{v}^{\ell,n}$ is detected as an anchored-bias vector across all datasets, and represents single space within the token because GPT2 tokeniser encodes same word with or without as different token numbers. For each value vector, we further unembeded the top-10 tokens, and most of them are human-interpretable words, which also verify that pretrained GPT2 family has intrinsic anchored bias within $W_{\text{out}}$.
Model Vector Top-10 Tokens
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) ␣The, ␣This, ␣A, ␣There, ␣It, ␣In, ␣We, ␣If, ␣When, ␣An
[…]
$\mathbf{v}^{9,788}$ (51.0%) ␣and, ␣in, ␣to, ␣or, ␣a, ␣at, ␣the, ␣..., ␣that, "
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) ␣a, ␣an, a, an, ␣another, ␣something, A, ␣the, ␣some, ␣any
$\mathbf{v}^{20,1731}(38.3\%)$first, ␣First, first, ␣FIRST, First, ␣1, 1, ␣Firstly, Firstly, ␣begin
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) ␣A, A, ␣An, ␣Aires, ␣Ae, ␣An, ierrez, AAF, Aim, ␣Aus
$\mathbf{v}^{34,2103}(70.2\%)$ romeda, maxwell, {CJK}UTF8gothドラゴン, lvl, mobi, elaide, amsung, nil, 911, dylib
$\mathbf{v}^{34,4178}$ (54.7%) reply, interstitial, ␣emer, 76561, ␣err, sg, ␣whe, eers, oi, ␣ignor
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) A, ␣A, a, AIN, aic, acebook, aa, An, AAAA, ae
$\mathbf{v}^{44,2995}$ (44.7%) ggles, atchewan, ␣aw, ␣let, eed, ␣ont, wn, ␣be, gg, ␣palp
$\mathbf{v}^{44,128}$ (11.9%) ␣A, ␣C, ␣E, ␣B, ␣S, ␣G, ␣F, ␣P, ␣D, ␣K
$\mathbf{v}^{38,4191}$ (100%) ␣a, ␣an, a, ␣of, ,, ␣and, ., ␣in, an, ␣the
$\mathbf{v}^{38,4174}$ (20.2%) ␣A, ,, ␣and, ␣a, ␣at, ., ␣of, ␣as, ␣on, ␣(
$\mathbf{v}^{37,423}$ (91.8%) ␣The, ␣This, ␣It, ␣There, ␣A, ␣If, ␣You, ␣We, ␣These, ␣When
$\mathbf{v}^{37,2966}$ (67.2%) ␣a, ␣an, ,, ␣and, ␣to, ␣the, ␣in, ␣(, ␣one, .
Figure 4: Attention pattern logit lens of the difference between anchored bias token ‘A’ and correct tokens (one of B,C,D,E), i.e., $\text{logit}_{T}^{\ell,h}[\texttt{A}](\mathbf{r}_{T}^{\ell,h})-\text{logit}_{T}^{\ell,h}[\texttt{B/C/D/E}](\mathbf{r}_{T}^{\ell,h})$ which is averaged within GPT-2 family across all layers and all datasets. The deeper the blue blocks are at each layer, the more serious the anchored bias is, and vice versa.
Locating attention heads of GPT2 family for anchored bias.
Following a similar method as locating anchored bias in MLP, we also aim to solve these research questions: 1) Is the attention head also responsible for the anchored bias in GPT2 family? 2) Which layer and head of attention pattern is anchored bias relevant to?
As explained in § 2, attention pattern $\mathbf{r}_{i,j}^{\ell,h}$ indicates the weighted average values where token $i$ attend to token $j$ by head $h$ at the layer $\ell$. We use logit lens to analyse the logit difference of final input prompt token contribution between anchored bias ‘A’ and correct choices ‘B/C/D/E’, i.e., $\text{logit}_{T}^{\ell,h}[\texttt{A}](\mathbf{r}_{T}^{\ell,h})-\text{logit}_{T}^{\ell,h}[\texttt{B/C/D/E}](\mathbf{r}_{T}^{\ell,h})$. As shown in Fig. 4, L8H1 and L10H8 in GPT2-Small, L18H12 and L20H5 in GPT2-Medium, L23H8 and L30H0 in GPT2-Large, L31H9 and L34H14666L34H14 indicates layer 34 and head 14 and both of them start from 0. in GPT2-XL are dominant heads related to anchored bias. Those heads are also distributed closer to the final layer in each GPT2 model. We zoom in on the L8H1 and L10H8 attention pattern of the final input token in GPT2-Small using a sample from IOI dataset. As shown in Fig. 5, the final token ‘:’ attends more weights on the anchored bias token ‘A’ than the correct choice token ‘B’, which agrees with our identified attention head using logit difference.
5 Mitigating Anchored Bias
Mitigating anchored bias in MLP.
Figure 5: The visualisation of identified anchored-bias attention head L8H1 and L10H8 in the GPT2-Small based on Fig. 4, where the attention weight of final token ‘:’ mainly attends to ‘A’ rather than ‘B’.
According to the findings in § 4, we localise the specific value vector in MLP related to the anchored bias. We further aim to solve the following research question: Can we fix the identified value vector in MLP by updating its values?
Following (Dai et al.,, 2022), we directly modify and update the identified value vector as:
[…]
Table 4: The new top-10 tokens of each updated value vector for each GPT2 model.
Model Updated Vector New Top-10 Tokens
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) ␣B, B, ␣b, ␣C, ␣D, ␣P, ␣L, ␣R, ␣H, ␣F
$\mathbf{v}^{9,2859}$ (61.8%) ␣B, B, ␣b, ␣C, ␣D, ␣L, ␣P, ␣R, ␣G, ␣F
$\mathbf{v}^{9,788}$ (51.0%) ␣B, ␣b, B, ␣C, ␣D, ␣L, ␣P, ␣R, ␣G, ␣BC
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) ␣C, C, ␣B, ␣c, ␣D, ␣G, ␣F, ␣P, ␣CS, ␣T
$\mathbf{v}^{20,1731}(38.3\%)$C, C, ␣B, ␣c, ␣D, ␣CS, ␣F, ␣P, ␣G, ␣T
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) ␣C, ␣A, ␣B, C, ␣c, ␣D, ␣F, ␣P, ␣G, ␣T
$\mathbf{v}^{34,2103}(70.2\%)$C, ␣c, C, ␣B, ␣D, ␣F, ␣P, ␣G, ␣T, ␣M
$\mathbf{v}^{34,4178}$ (54.7%) ␣C, ␣c, C, ␣B, ␣P, ␣D, ␣F, ␣G, ␣T, ␣L
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) ␣C, ␣c, C, ␣A, ␣B, ␣D, ␣F, ␣P, ␣T, ␣G
$\mathbf{v}^{44,2995}$ (44.7%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣T, ␣F, ␣R, ␣G
$\mathbf{v}^{44,128}$ (11.9%) ␣A, ␣C, ␣E, ␣B, ␣S, ␣G, ␣F, ␣P, ␣D, ␣K
$\mathbf{v}^{38,4191}$ (100%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣F, ␣T, ␣L, ␣R
$\mathbf{v}^{38,4174}$ (20.2%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣F, ␣T, ␣L, ␣S
$\mathbf{v}^{37,423}$ (91.8%) ␣C, ␣c, C, ␣B, ␣D, ␣F, ␣P, ␣T, ␣L, ␣G
$\mathbf{v}^{37,2966}$ (67.2%) ␣C, ␣c, C, ␣B, ␣D, ␣P, ␣F, ␣T, ␣G, ␣L
where $\lambda_{1}=1,\lambda_{2}=8$. After updating the corresponding value vector in MLP, we utilise the updated GPT2 model to predict the next token of the same input MCQ prompts. We comprehensively evaluate each updated GPT2 model with the corresponding modified value vector using Infer. and Eva. across all datasets. As shown in Table 5, most updated value vectors achieve high classification accuracy regarding MCQ tasks, even with multiple near 100% or 100% accuracy in $\mathbf{v}^{9,1853}$ and $\mathbf{v}^{9,788}$ of GPT2-Small, $\mathbf{v}^{34,1541}$ of GPT2-Large, $\mathbf{v}^{44,4967}$ and $\mathbf{v}^{38,4191}$ of GPT2-XL. For those value vectors with around a 60-70% chance of anchored bias happening across all datasets and different GPT2 models (i.e., $\mathbf{v}^{9,2859}$ of GPT2-Small, $\mathbf{v}^{20,3713}$ of GPT2-Medium, $\mathbf{v}^{34,2103}$ of GPT2-Large, and $\mathbf{v}^{37,2966}$ of GPT2-XL), the classification accuracy still 68.09% averaged all datasets and all models. This means that the simple and straightforward method (i.e., Eq. 10) is effective, and we do not need to fine-tune the whole GPT2 model to fix the anchored bias. We further unembeded the updated value vectors in Table 4, and it shows that the anchored bias token ‘A’ is significantly removed and the new top-10 tokens for each new value vector are replaced with the correct choices token ‘B/C/D/E’.
Table 5: The classification accuracy of each MCQ inference (Infer.) and evaluation (Eva.) dataset after updating the identified value vectors for each GPT2 model.
Model Vector IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small $\mathbf{v}^{9,1853}$ (100%) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
$\mathbf{v}^{9,2859}$ (61.8%) 44.6 48.9 100.0 100.0 86.5 59.4 95.3 95.9 59.4 58.8
$\mathbf{v}^{9,788}$ (51.0%) 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0
GPT2-Medium $\mathbf{v}^{20,3713}$ (79.3%) 99.4 100.0 97.1 100.0 68.3 25.3 61.9 61.4 21.9 21.3
$\mathbf{v}^{20,1731}(38.3\%)$ 33.6 36.1 80.0 55.6 0.0 0.0 54.3 50.6 20.6 22.5
GPT2-Large $\mathbf{v}^{34,1541}$ (100%) 100.0 100.0 100.0 100.0 100.0 100.0 98.2 100.0 99.5 99.0
$\mathbf{v}^{34,2103}(70.2\%)$ 71.2 75.0 55.0 65.0 94.4 32.3 51.7 59.1 46.8 63.9
$\mathbf{v}^{34,4178}$ (54.7%) 25.1 26.0 8.9 15.0 93.8 35.4 47.4 48.9 49.7 62.9
GPT2-XL $\mathbf{v}^{44,4967}$ (98.0%) 98.2 98.8 100.0 100.0 100.0 100.0 93.9 77.8 95.7 100.0
$\mathbf{v}^{44,2995}$ (44.7%) 24.8 38.8 100.0 100.0 100.0 99.0 96.8 81.0 95.3 96.8
$\mathbf{v}^{44,128}$ (11.9%) 67.8 72.9 96.8 100.0 100.0 100.0 94.9 79.4 94.6 96.8
$\mathbf{v}^{38,4191}$ (100%) 100.0 100.0 100.0 100.0 100.0 94.9 99.6 88.9 97.9 100.0
$\mathbf{v}^{38,4174}$ (20.2%) 11.4 7.1 32.3 33.3 68.5 51.0 22.2 22.2 9.1 12.6
$\mathbf{v}^{37,423}$ (91.8%) 54.3 55.3 100.0 100.0 100.0 92.9 83.0 57.1 41.8 49.5
$\mathbf{v}^{37,2966}$ (67.2%) 34.0 40.0 100.0 100.0 98.4 81.6 79.3 57.1 56.7 56.8
Mitigating anchored bias in attention heads.
Based on the located attention heads for each GPT2 model in § 4, we follow the same pattern as fixing anchored bias in MLP and further propose a recalibration approach to mitigate the anchored bias in the attention head by swapping the attention weight of $\mathbf{r}_{T}^{\ell,h}$ between the position of ‘A’ and ‘B/C/D/E’:
$\mathbf{r}_{T,p(\texttt{A})}^{\ell,h}=\mathbf{r}_{T,p(\texttt{B/C/D/E})}^{\ell,h}\quad\mathbf{r}_{T,p(\texttt{B/C/D/E})}^{\ell,h}=\mathbf{r}_{T,p(\texttt{A})}^{\ell,h}$ (11)
[…]
Table 6: The Top-1 classification accuracy changes after attention pattern recalibration for each GPT2 model across all datasets.
Model Head IOI (2) LD (3) Greater (4) ARC (4) CSQA (5)
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
GPT2-Small L8H1 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.04 0.0 0.0
L10H8 0.0 0.0 0.0 0.0 0.0 0.0 0.0 2.04 0.0 0.0
GPT2-Medium L18H12 92.47 90.72 21.76 27.78 3.39 3.16 24.33 21.69 1.67 1.25
L20H5 0.34 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
GPT2-Large L23H8 0.0 0.0 0.0 0.0 0.0 0.0 0.38 1.14 0.0 2.06
L30H0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
GPT2-XL L31H9 0.0 0.0 9.68 0.0 0.57 0.0 0.53 0.0 0.12 0.0
L34H14 0.0 0.0 12.90 0.0 0.57 0.0 3.33 1.67 0.46 0.0
6 Discussion
Is Few-shot learning helpful?
Based on the comprehensive zero-shot learning MCQ across all datasets, we have a good understanding of how important the MLP and attention head are regarding the anchored bias in the GPT2 family. The following question might be whether few-shot learning can mitigate this anchored bias without updating specific value vectors in MLP or recalibrating the attention head. We conduct a preliminary experiment to evaluate GPT2 family across all datasets using 1-shot and 2-shot learning settings. The initial finding is that the anchored bias could be relatively mitigated to a certain extent depending on the length of few-shot prompts and the complexity of the few-shot prompt content. However, the anchored bias still happens, especially for the long and complex few-shot prompts. Even if the anchored bias to ‘A’ disappears, GPT2 model cannot still predict the correct choice and generate other incorrect choices, which needs more investigation. We leave this question as the future work.
Is direct value vector updating in MLP harmful to the ability of GPT2 for other tasks?
We conduct a preliminary experiment to evaluate whether direct value vector updating in MLP is harmful to the ability of GPT2-Small for the original IOI task. The initial result is that modified GPT2-Small can still achieve 85.6% accuracy to successfully predict the non-subject of the last clause. For a better and minimal-harm MLP updating algorithm development, we leave it for future work.
7 Related Work
Mechanistic Interpretability.
Our work falls into the scope of investigating Mechanistic Interpretability, a field studying neural networks with an intent to understand and explain their inner workings. Recent studies have advanced our understanding of how factual associations are managed within transformer models. Geva et al., (2023) investigates the information flow during inference to understand how facts are internally retrieved. Meng et al., 2022a focuses on the storage of knowledge in GPT-like autoregressive transformer models, noting significant differences from masked models, such as unidirectional attention and generation capabilities. They use causal mediation analysis to examine how hidden state activations influence fact recall and introduce a novel model editing method that alters model behaviour, providing deeper insights into the neural mechanisms of information retrieval. Building on these findings, Meng et al., 2022b develops a scalable multi-layer update algorithm that targets transformer weights to enhance memory recall. Similarly, Gupta et al., (2023) proposes a model editing strategy to modify commonsense knowledge in GPT-2 models. Halawi et al., (2023) identifies certain layers of language models following incorrect demonstrations that lead to harmful behaviours. They utilise logit lens that analyses model internal behaviour. Wang et al., (2022) concentrates the task of IOI, and identifies critical circuits of 26 attention heads. Our work also focuses on the task of IOI and reformulates the task as a binary question-answering task to investigate LLM’s positional bias.
Positional Bias of LLMs in MCQ.
Various studies have highlighted the impact of positional bias on the accuracy of LLMs when addressing multiple-choice questions. Pezeshkpour and Hruschka, (2023) demonstrate a significant "sensitivity gap" in LLMs, such as GPT-4, attributing a drop in performance of up to 75% in the zero-shot setting to positional bias. They propose calibration strategies that have successfully improved accuracy by up to 8 percentage points. Similarly, Wang et al., (2023) identify a notable influence of display order on GPT-4’s evaluation scores, advocating a calibration framework that includes Multiple Evidence Calibration, Balanced Position Calibration, and Human In The Loop Calibration. This approach has notably increased evaluation accuracy and better aligned LLM outputs with human judgments. Zheng et al., 2024a explore the "selection bias" of LLMs, where models prefer certain option labels such as "Option A" more frequently. Their novel debiasing technique, PriDe, effectively mitigates this bias by adjusting the prediction distribution at inference time without the need for labels. Wang et al., (2024) further investigate how altering the order of answer options affects LLM understanding, observing significant performance degradation when answer options are reordered. This body of work collectively underscores the necessity for robust strategies to address and correct positional biases in LLMs, ensuring more reliable model performance in multiple-choice settings. Turpin et al., (2024) explores the impact of positional bias of MCQA for Chain-of-thought (CoT) generation. The key findings reveal that the introduction of biasing features significantly affects the CoT predictions. When biasing features are applied, models modify their explanations to support incorrect, bias-consistent answers, occasionally providing sound yet unfaithful reasoning. This suggests need for robust mechanisms to detect and mitigate biases. Zheng et al., 2024b investigate the application of advanced LLMs in the role of evaluators, particularly for assessing responses to open-ended questions. They delve into the strengths and weaknesses of employing LLMs as judges, highlighting biases related to position, verbosity, self-enhancement biases, and constrained logical capabilities. Furthermore, they suggest strategies to alleviate some of these issues.
8 Conclusion
In this work, we identify the anchored bias of GPT2 family, where GPT-2 models consistently favour the first choice ‘A’ in the MCQ task. Based on this observation, we comprehensively conduct a mechanistic analysis of the internal workings of GPT2 family. We find that some value vectors in MLP modules with specific layers and dimensionality play a significant role in the anchored bias, and we further use a straightforward but potent approach to update the corresponding value vectors, which effectively mitigate the anchored bias in GPT2 family. In addition, some attention heads also play auxiliary roles in this bias, and the recalibration approach works well for the IOI dataset in GPT2-Medium.
Appendix A The Statistic Information of Each Test Dataset for GPT2 Models
We split each dataset into 90% Infer. set for anchored bias discovering and mitigation, and 10% Eva. set for modified GPT2 model verification on the MCQ task.
Table 7: Statistic of the test datasets for each GPT2 model, where Infer. represents 90% test dataset used to discover and mitigate anchored bias, and Eva. represents 10% test dataset used to verify the performance of updated GPT2 models.
Num. GPT2-Small GPT2-Medium GPT2-Large GPT2-XL
Infer. Eva. Infer. Eva. Infer. Eva. Infer. Eva.
IOI (2) 410 45 877 97 900 100 773 85
LD (3) 114 12 170 18 180 20 31 3
Greater (4) 289 32 856 95 897 99 883 98
ARC (4) 446 49 748 83 797 88 571 63
CSQA (5) 308 34 720 80 881 97 864 95